Excrcise 16: Relation of rights to EPS and the price-carnings ratio Walker Machine Tools has 7 million shares of common stockjust
Answer #1

ANSWER (a)

Before the New Shares issue

Earning Per Share (EPS)= Net Income/ No. of Shares

$2,50,00,00/70,00,000= 3.57

Price Earning Ratio= Stock Price/EPS

82/3.57=22.97

b. After right offering

calculation of theoretical ex right fair value per share

(Fair Value of Shares before issue of right shares+ total amount received on right)/(No. of shares before right share issue+ No. of shares to be issued)

=(70,00,000*82)+(7,00,000*76.50)/70,00,000+7,00,000

=$81.50 per share

Computation of Adjustment Factor=

  Fair Value per share before issue of right shares/theoretical ex right fair value per share

=82/81.50=1.006

Earning Per Share (EPS) = Net Income/ No. of Shares

(after issue of Right Share) $2,50,00,00/70,00,000/1.006= 3.57/1.006=3.55

Price Earning Ratio = Stock Price/EPS

(after issue of Right Share) 82/3.55=23.09